# WildAvatar: Web-scale In-the-wild Video Dataset for 3D Avatar Creation

### v2

@@ Abstract
Existing human datasets for avatar creation are typically limited to laboratory environments, wherein high-quality annotations (e.g., SMPL estimation from 3D scans or multi-view images) can be ideally provided. However, their annotating requirements are impractical for real-world images or videos, posing challenges toward real-world applications on current avatar creation methods. To this end, we propose the WildAvatar dataset, a web-scale in-the-wild human avatar creation dataset extracted from YouTube, with $10,000+$ different human subjects and scenes. WildAvatar is at least 10×10$\times$10 × richer than previous datasets for 3D human avatar creation. We evaluate several state-of-the-art avatar creation methods on our dataset, highlighting the unexplored challenges in real-world applications on avatar creation. We also demonstrate the potential for generalizability of avatar creation methods, when provided with data at scale. We publicly release our data source links and annotations, to push forward 3D human avatar creation and other related fields for real-world applications.
Figure 1: Examples from our WildAvatar dataset. Unlike previous laboratory datasets for 3D avatar creation, WildAvatar curates in-the-wild web videos. With $10,000+$ human subjects and scenes, WildAvatar is at least 10×10$\times$10 × richer than the previous datasets, covering a wide range of real-world human motions and appearances. It contains high-quality annotations (i.e., SMPL, camera parameters, and human segmentation masks).
@@ 1 Introduction
FreeMan [69] 40/123 ITW [Uncaptioned image]+[Uncaptioned image] SynWild [20] 5/5 ITW --
NeuMan [31] 6/6 ITW -- TikTok [30] 340/340 ITW --
IVS-Net [15] <700/<700 ITW -- WildAvatar 10k+/10k+ ITW --
@@
3D Human Avatar creation has extensive applications in VR/AR, film-making, metaverse, etc., attracting significant attention recently. With the advent of neural radiance fields (NeRF) [49, 66], recent research aims to recover 3D avatars from 2D observations, enabling the synthesis of novel images from arbitrary viewpoints and body poses [72, 53, 51, 26, 18, 38, 25]. For example, many works focus on reconstructing animatable human models from well-annotated and calibrated multi-view images and videos [53, 51, 52, 73], and have achieved photo-realistic results in laboratory datasets [53, 56, 4]. However, the well-designed collection systems limit its scalability to real-world scenes. To mitigate the gap between the ideal laboratory setting and real-world scenarios, a line of research proposes to reconstruct human avatars from monocular videos [19, 26, 72] or even a single image [25, 15], considering these forms of data can be easily accessible in our daily life. Still, these methods require annotations from multi-view images or even 3D scans. In addition, their performances are restricted by small-scale human data, especially for generalizable creation.
@@
Although previous arts attempted to make existing laboratory datasets more suitable for real-world applications, there still exists a domain gap between laboratory and real-world scenarios. This is down to the fact that current human datasets are mainly collected via well-designed laboratory systems, which are expensive and time-consuming to collect. As shown in Tab. 1, existing human data collections mostly rely on accurate annotations from various advanced devices, such as well-calibrated multi-view cameras [78, 63, 48, 14], depth sensors [9, 84, 78, 28], IMUs [68], or expensive scanners [56, 76], as well as specialized actors and lightstages [21, 77, 67, 69, 42]. This ideal condition is unavailable in in-the-wild scenarios (i.e., monocular web videos) or consumer applications. To tackle this problem, recent efforts are attempting to collect in-the-wild monocular human data [30, 15, 17] but rely heavily on costly manual interventions, making it difficult to scale up. Therefore, their inadequate diversity fails to meet the requirements for dealing with in-the-wild challenges of 3D avatar creation. Given the shortage of in-the-wild human data, scaling up avatar creation from real-world scenarios is worth exploring.
@@
To facilitate the in-the-wild avatar creation, we curate WildAvatar, a large-scale in-the-wild human avatar creation dataset extracted from YouTube, with $10,000+$ different human subjects and scenes. Specifically, we collect monocular human videos from YouTube and streamline the annotation process with off-the-shelf state-of-the-art annotation methods [55, 36, 2, 34], e.g., YOLO [55], Segment Anything [34]. Without costly sensors or lightstages, we collect over $10,000$ in-the-wild human samples. WildAvatar fills the gap in large-scale in-the-wild human data collection, offering at least 10×10$\times$10 × richer human subjects than the existing human datasets.
@@
To investigate the performance of existing avatar creation methods in real-world scenarios, we evaluate state-of-the-art avatar creation methods [53, 51, 72, 26] on our WildAvatar dataset. Compared to laboratory datasets, existing methods encounter a significant performance drop in the reconstruction quality of the WildAvatar dataset. This highlights the gap between well-annotated laboratory datasets and in-the-wild datasets for avatar creation. We also investigate the scalability of existing generalizable human avatar creation methods [18, 38, 25] with WildAvatar, which suggests a significant improvement (up to $7\%$) in real-world scenarios. We will release the WildAvatar dataset, providing the video IDs, frame IDs, extracting scripts, and annotations obtained from our pipeline. We hope that our dataset will push forward the development of 3D human avatar creation, and other related topics, e.g., human mesh estimation (HPS) [33, 36, 41, 79, 37, 35, 27] and 3D human avatar generation [10, 23, 12, 24, 70, 46]. Overall, our main contributions are summarized as follows:
@@
  • We present WildAvatar, a large-scale avatar dataset collected from in-the-wild videos with more than $10,000$ human subjects, at least 10×10$\times$10 × larger than previous datasets. The scale-up facilitates the creation of per-subject human avatars and paves a new avenue for generalizable human avatar reconstruction.
  • We evaluate the performance of existing state-of-the-art human avatar creation methods on the collected WildAvatar, showing the critical challenges of the in-the-wild avatar creation. Our observations could inspire the key directions in the human avatar creation.
@@
  • We will release our data collection and annotation pipeline for WildAvatar. Further data scale-up for large-scale model training will be unlocked. The open data and code could provide important insights for future avatar creation and relevant human generation models.
@@ 2 Related Work
3D Avatar Creation Datasets. Tab. 1 presents a system overview of previous avatar creation datasets. Previous datasets mainly obtain high-quality human mask and SMPL annotations with ideal laboratory systems, e.g., well-calibrated multi-view cameras [53, 9, 82, 78, 28, 63, 83, 62, 76, 14, 21, 81, 85, 29, 77, 73, 67, 48, 42, 68], depth sensors [9, 84, 78, 28, 63, 14], IMUs [68], or expensive scanners [56, 76], as well as specialized actors and lightstages [53, 56, 9, 84, 82, 78, 28, 63, 83, 62, 76, 14, 21, 81, 85, 29, 77, 73, 67, 48, 42, 68], depth sensors [9, 84, 78, 28, 63, 14]. Unfortunately, in real-world scenarios or applications, these optimal conditions are typically hard to guarantee. Towards in-the-wild avatar creation, previous efforts also attempt to collect web human videos [30, 15]. However, they rely heavily on costly manual interventions (e.g., pre-screening or mask extraction based for  [44]) and fail to scale up. And they either show little viewpoint change [30] or not public released [15]. Therefore, there are still significant demands for a public large-scale in-the-wild dataset for avatar creation.
3D Avatar Creation Methods. Existing works on avatar creation learn coarse human surfaces with parametric mesh models [6, 47, 50, 74, 57], but these explicit meshes cannot express detailed geometry or appearance. Pifu-based methods [58, 59, 22, 40, 16] represent the human body with pixel-aligned functions and require high-quality synthetic data, but fail to generalize to real-world scenarios due to the domain gap. NeRF-based approaches [53, 51, 72, 52, 19, 39, 8, 75] learn implicit human representation from high-quality videos, and achieve photo-realistic rendering in the laboratory benchmarks. In addition, generalizable models [18, 38, 25] further simplify the data demand to a single image. Nonetheless, these methods still rely on accurate annotations, which only exist in ideal laboratory environments. Recent attempts [31, 20] decompose avatars and scenes for construction at once, demanding accurate global alignments between human bodies and backgrounds that are not available in real-world videos.
@@
SMPL Estimation in the Wild. Parametric body models (e.g., SMPL [47], SMPLX [50]) encode coarse human surfaces with pose and shape parameters for avatar creation. For example, end-to-end methods [33, 36, 79, 41, 27] estimate the SMPL parameters efficiently from a single in-the-wild image, but they are not robust with complex scenes. Subsequent works [37, 32] refine end-to-end estimation results via fitting SMPL to 2D annotations in the loop. Recent works further consider imposing temporal consistency [2, 65, 35] to smooth SMPL estimations across input in-the-wild videos. Yet, it is far from a comprehensive annotating streamline for real-world video sources.

### v4

# WildAvatar: Learning In-the-wild 3D Avatars from the Web

@@ Abstract
Existing research on avatar creation is typically limited to laboratory datasets, which require high costs against scalability and exhibit insufficient representation of the real world. On the other hand, the web abounds with off-the-shelf real-world human videos, but these videos vary in quality and require accurate annotations for avatar creation. To this end, we propose an automatic annotating pipeline with filtering protocols to curate these humans from the web. Our pipeline surpasses state-of-the-art methods on the EMDB benchmark, and the filtering protocols boost verification metrics on web videos. We then curate WildAvatar, a web-scale in-the-wild human avatar creation dataset extracted from YouTube, with $10,000+$ different human subjects and scenes. WildAvatar is at least 10×10$\times$10 × richer than previous datasets for 3D human avatar creation and closer to the real world. To explore its potential, we demonstrate the quality and generalizability of avatar creation methods on WildAvatar. We will publicly release our code, data source links and annotations to push forward 3D human avatar creation and other related fields for real-world applications.
Figure 1: Overview of WildAvatar. (a) Unlike previous laboratory datasets for 3D avatar creation, WildAvatar curates in-the-wild web videos. (b) With $10k+$ human subjects and scenes, WildAvatar is at least 10×10$\times$10 × richer than the previous datasets. (c) It contains high-quality annotations and demonstrates impressive potential to boost the quality and generalizability of avatar-creation methods.
$\dagger$: Corresponding author.
@@ 1 Introduction
FreeMan [80] 40/123 ITW [Uncaptioned image]+[Uncaptioned image] SynWild [25] 5/5 ITW --- -
NeuMan [37] 6/6 ITW --- - TikTok [36] 340/340 ITW --- -
IVS-Net [18] <700/<700 ITW --- - WildAvatar 10k+/10k+ ITW --- -
@@
3D Human Avatar creation has extensive applications in VR/AR, film-making, metaverse, etc., attracting significant attention recently. With the advent of neural radiance fields (NeRF) [57, 77], recent research aims to recover 3D avatars from 2D observations, enabling the synthesis of novel images from arbitrary viewpoints and body poses [84, 60, 59, 29, 21, 45, 31]. For example, many works focus on reconstructing animatable human models from well-annotated and calibrated multi-view images and videos [60, 59, 61, 85], and have achieved photo-realistic results in laboratory datasets [60, 64, 4]. However, their performances are restricted by small-scale indoor human data, especially for generalizable creation. This is down to the fact that current avatar creation datasets are mainly collected via well-designed laboratory systems, which are expensive and time-consuming to collect. In addition, there still exists domain gaps between laboratory and real-world scenarios. As shown in Table 1, existing human data collections mostly rely on annotations from advanced devices, such as well-calibrated multi-view cameras [95, 73, 56, 16], depth sensors [9, 100, 95, 33], IMUs [79], or expensive scanners [64, 93], as well as specialized actors and lightstages [26, 94, 78, 80, 50]. These ideal conditions require high costs against scalability, exhibit limited representations of the real world, and are unavailable in in-the-wild scenarios (i.e., monocular web videos) or consumer applications.
@@
To tackle this problem, recent efforts are attempting to collect in-the-wild monocular human data [36, 18, 20] from the web. However, they still rely heavily on costly manual interventions, making it difficult to scale up. Therefore, their inadequate diversity fails to meet the requirements for dealing with in-the-wild challenges of 3D avatar creation. Given the shortage of real-world human data, scaling up avatar creation from real-world scenarios is worth exploring.
To this end, we propose a novel pipeline with filtering protocols to extract human movements automatically from the web. Specifically, we first streamline the annotation process with off-the-shelf state-of-the-art annotation methods [63, 23, 2, 41], e.g., YOLO [63], Segment Anything [41]. We then propose a suite of assessment protocols for automatic filtering to retain only qualified video clips (e.g., without occlusions and annotated in high confidence). We report our performance on the EMDB benchmark to investigate our pipeline’s performance in real-world scenarios. With the help of the additional stage for aligning and smoothing, our pipeline surpasses state-of-the-art methods (up to $4\%$). We also demonstrate that the proposed filtering protocols can boost verification metrics among web video sources.
@@
We implement the proposed annotating and filtering pipeline to collect real-world human samples from the web. Without costly sensors or lightstages, we curate WildAvatar, a large-scale in-the-wild human avatar creation dataset extracted from YouTube, with $10,000+$ different real-world human subjects and scenes. WildAvatar fills the gap in large-scale in-the-wild human data collection, offering at least 10×10$\times$10 × richer human subjects than the existing human datasets.
@@
To explore the potential of WildAvatar, we investigate the scalability of existing generalizable human avatar creation methods [21, 45, 31] with WildAvatar, which suggests a significant improvement (up to $7\%$) in real-world scenarios. We will release the WildAvatar dataset, providing the video IDs, frame IDs, extracting scripts, and annotations obtained from our pipeline. We hope that our dataset will push forward the development of 3D human avatar creation and other related topics, e.g., human mesh estimation (HPS) [39, 43, 48, 96, 44, 42, 32, 71], 3D human avatar generation [11, 28, 14, 30, 81, 54], 3D human interactions [82, 72], and 3D relightable avatar [88, 34, 13]. Overall, our main contributions can be summarized as follows:
  • We propose a novel annotating pipeline with filtering protocols to curate human movements from the web. Our pipeline surpasses state-of-the-art methods on the in-the-wild benchmark, and our filtering protocols can boost verification metrics on annotating web video sources.
@@
  • We curate WildAvatar, a large-scale avatar dataset collected from in-the-wild videos with $10,000+$ human subjects, at least 10×10$\times$10 × larger than previous datasets. The scale-up facilitates the creation of per-subject avatars and paves a new avenue for generalizable avatar reconstruction.
@@
  • We illustrate the great potential of our pipeline and WildAvatar, with exploratory experiments on supporting popular downstream 3D avatar creation applications. Further data scale-up for large-scale model training will be unlocked. The open code and data could provide important insights for future avatar creation and relevant tasks.
@@ 2 Related Work
SMPL Annotation in the Wild. Parametric models (e.g., SMPL [55], SMPLX [58]) encode coarse human surfaces with pose and shape parameters for 3D avatar creation. For example, end-to-end methods [39, 43, 96, 48, 32] estimate the SMPL parameters efficiently from single in-the-wild images, but they are not robust with complex scenes. Subsequent works [44, 38] refine end-to-end estimation results via fitting SMPL to 2D annotations in the loop. Recent works further consider imposing temporal consistency [2, 75, 42] to smooth SMPL estimations across in-the-wild videos. Yet, it is far from achieving a comprehensive annotating and filtering streamline for real-world video sources.
3D Avatar Creation Datasets. Table 1 presents a system overview of previous avatar creation datasets. Previous datasets mainly obtain high-quality human mask and SMPL annotations with ideal laboratory systems, e.g., well-calibrated multi-view cameras [60, 9, 99, 95, 33, 73, 101, 70, 93, 16, 26, 98, 102, 35, 94, 85, 78, 56, 50, 79], depth sensors [9, 100, 95, 33, 73, 16], IMUs [79], or expensive scanners [64, 93], as well as specialized actors and lightstages [60, 64, 9, 100, 99, 95, 33, 73, 101, 70, 93, 16, 26, 98, 102, 35, 94, 85, 78, 56, 50, 79], depth sensors [9, 100, 95, 33, 73, 16]. Unfortunately, these optimal conditions are typically high-cost against scaling up and limited in scenarios compared to the real world. Towards in-the-wild avatar creation, previous efforts also attempt to collect avatar data from web human movement videos [36, 18]. However, they rely heavily on costly manual interventions (e.g., pre-filtering or mask extraction based on  [52]) and still fail to scale up. And they either show little viewpoint change [36] or not public released [18]. Therefore, there are still significant demands for a public large-scale in-the-wild dataset for avatar creation.
3D Avatar Creation Methods. Existing works on avatar creation learn coarse human surfaces with parametric mesh models [6, 55, 58, 87, 65], but these explicit meshes cannot express detailed geometry or appearance. Pifu-based methods [66, 67, 27, 49, 19] represent the human body with pixel-aligned functions and require high-quality synthetic data, but fail to generalize to real-world scenarios due to the domain gap. NeRF-based approaches [60, 59, 84, 61, 22, 47, 8, 91] learn implicit human representation from high-quality videos, and achieve photo-realistic rendering in the laboratory benchmarks. In addition, generalizable models [21, 45, 31] further simplify the data demand to a single image. Nonetheless, these methods still rely on accurate annotations, which only exist in ideal laboratory environments. Recent attempts [37, 25] decompose avatars and scenes for construction at once, demanding accurate global alignments between human bodies and backgrounds that are not available in real-world videos. Web-scale in-the-wild video data is crucial for the next phase of 3D avatar development.
3 Methodology
Our goal is to build a large-scale in-the-wild dataset for human avatar creation. To this end, we collect abundant real-world human videos from YouTube and annotate corresponding labels. Without time-consuming and high-cost manual filtering and annotation [18], we design an efficient pipeline to filter video candidates and obtain high-quality annotations. Specifically, we download human movement videos from the web with our automatic tools, as described in Fig. 2. Then, we collect high-quality annotated data with several processing steps on downloaded web videos (e.g., annotating in Section 3.2 and filtering in Section 3.3). Finally, we evaluate our designs in detail in Section 3.4.
3.1 Data Collection
Figure 2: The four-stage data processing pipeline. We first obtain the bounding box of key subjects in videos in Stage I and extract human segmentation masks in Stage II. Then, the SMPL and camera parameters are coarsely estimated in Stage III and later refined in Stage IV.
We follow two steps to collect human movement videos: 1) We first search for a large and diverse set of video candidates on YouTube. Note that video candidates could contain some noisy videos without human subjects; 2) We then download the candidate videos with automatic tools [3], cut them into manageable clips, and filter out short video clips. Details can be found in Appendix B.
3.2 Data Annotating
To obtain high-quality annotations (i.e., SMPL, camera parameters, and human segmentation mask) for the collected human videos, we design the following four stages to annotate the in-the-wild human video clips automatically.
Stage I: Human Bounding Box Detection and Tracking. Precise human bounding boxes are important for monocular SMPL estimation methods [43, 39, 48, 96] and human segmentation (using Segment Anything [41]). Hence, we first obtain the bounding box of human subjects with off-the-shelf state-of-the-art detection methods [63, 86].
Stage III & IV: SMPL and Camera Estimation. We first estimate SMPL and camera parameters frame by frame, using state-of-the-art single-image-based human pose and shape estimation methods [43, 23] in Stage III. However, these coarse parameters may ignore the temporal consistency of human motions. Inspired by previous SMPL annotation pipelines [44, 2, 38], we further refine SMPL and camera parameters in Stage IV, smoothing these parameters across the whole video clip sequence via gradient descent. We also incorporate the estimated 2D keypoints and SAM masks into a refinement loop, providing additional supervision for precise SMPL and camera parameter estimation. Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in Fig. E.
3.3 Data Filtering
Figure 3: Visualizations of filtering protocols. We only retain video clips that (a) show high confidence in human detections; (b) obtain high average confidence in 2D pose estimations; (c) consistency annotated by different expert models; (d) consistency on keypoints between projected SMPL keypoints and 2D pose estimations; and (e) consistency on segmentation masks between Segment-Anything and SMPL.
In contrast to high-quality human videos collected in well-designed laboratory systems, real-world videos are not always qualified for annotating or curating into datasets. To automatically filter out these unqualified videos (e.g., in severe occlusions or low resolution), we propose a suite of assessment protocols within the four-stage pipeline.
Protocol I: Clear Body with Significant Movement. The primary focus is to ensure the presence of a clear human body with no occlusion. To this end, we only retain in-the-wild videos with high confidence in both detection and 2D pose estimation, as illustrated in Fig. 3 (a) and (b). Then, we further reduce occlusion by only focusing on the key subject within each video clip. This is achieved by calculating the average size of the bounding box that encloses the tracked subject throughout the entire video. We also eliminate those trivial human movement videos, specifically those too brief or with insignificant viewpoint shifts and human movements.
Protocol II: Ensemble of Annotating Experts. To ensure the quality of in-the-wild annotation, we adopt the ensemble of different state-of-the-art annotating model experts. Specifically, we denote the model libraries for detection [63, 86], 2D pose estimation [74, 92, 10], and SMPL Estimation [43, 23] as $\{\mathcal{D},\mathcal{J},\mathcal{S}\}$, and their predictions as $\{D_{i},J_{i},S_{i}\}$. We calculate the average $\{\mu(D_{i}),\mu(J_{i}),\mu(S_{i})\}$ as the fused predictions, and the standard deviation $\{\sigma(D_{i}),\sigma(J_{i}),\sigma(S_{i})\}$ as references for filtering. To reduce annotating errors, we only retain the videos that are annotated consistently among different state-of-the-art annotating experts (see Fig. 3 (c)).
Protocol III: Consistency on 2D Keypoints. To augment the reliability of SMPL estimations, we implement an additional double-check of the consistency in 2D keypoints, between the monocular SMPL estimation and 2D pose estimation results. Specially, we project the 3D SMPL keypoints $J_{3D}(\mu(S_{i}))$ to 2D. Then, we compute the Percentage of Correct Keypoints (PCK) [90] between the projected $\mathbf{\Pi}(J_{3D}(\mu(S_{i})))$ and those predicted from 2D pose estimation $\mu(J_{i})$. Subsequently, we discard video clips that exhibit low average PCK values to ensure that only videos with high annotation confidence are retained for further curation.
Protocol IV: Consistency in SMPL and SAM Masks. With inaccurate predictions, monocular SMPL estimation and SAM might have low overlaps in estimated human masks. We double-check their consistency by comparing SMPL projection masks with the SAM masks. Intuitively, the SMPL mask denotes the naked body, while the SAM mask contains the clothed body. Therefore, the SMPL mask should be mostly covered by the SAM mask. To guarantee high-quality annotations, we discard video clips whose SAM-masked region has small overlaps with the SMPL-masked region, as illustrated in Fig. 3 (e) (more examples in Fig. F).
3.4 Analysis
Evaluation of Annotating Pipeline. To demonstrate the accuracy of our four-stage pipeline among in-the-wild videos, we first provide quantitative analysis on real-world EMDB [40] benchmark in Table 2. Due to the additional 2D aligning and temporal smoothing in Stage IV, our annotations have more accurate alignment on extremities (e.g., feet, arms, and heads). Our pipeline surpasses existing methods, especially on the MPJPE [33] ($+3.27\%$) and PVE [89] ($+4.16\%$), where global alignments are not taken into account. Evaluation details and qualitative comparisons can be found in Appendix C.
[…]
SPIN [44] 87.1 140.3 174.9
VIBE [42] 81.4 125.9 146.8
PARE [43] 72.2 113.9 133.2
TRACE [76] 70.9 109.9 127.4
CLIFF [48] 68.1 103.3 128.0
TCMR [17] 65.6 119.1 137.7
HybrIK [46] 65.6 103.0 122.2
HMR2.0 [23] 60.6 98.0 120.3
Ours 59.9 94.9 115.5
Table 2: Performance comparisons on the public EMDB [40] benchmark. Our proposed four-stage pipeline surpasses existing methods on the quality of SMPL annotations.
Verification of Filtering Strategies. We further demonstrate the efficacy of our filtering protocols and validate the quality of our final dataset. Given that there is no ground truth (GT) for evaluation, we adopt double-checks for verification to assess the quality of annotations for in-the-wild data. Specifically, we evaluate the consistency between predicted annotations after applying our protocols. The detailed results are presented in Table 3, where PCK represents the consistency between the 2D pose estimation models and the SMPL models, while the SOIOU measures the proportion of the SMPL masks that lie outside the SAM masks. Comparing adjacent columns, we find our pipeline designs compelling: 1) Protocol I is a vital pre-processing, which brings a marginal improvement on PCK and SOIOU. It eliminates a substantial number of clips in poor quality (e.g., no human subjects or movements); 2) Protocol II also yields a significant enhancement on PCK ($10.1\%$), demonstrating the importance of adopting various annotating experts. 3) Protocol III & IV step further towards consistency, raise PCK by $7.51\%$ and drop SOIOU by $0.094$ in total; and 4) Stage IV finally refines the SMPL annotations, resulting in PCK over $0.92$ and SOIOU lower than $0.03$. Notably, this final PCK deviates by only $1.7\%$ from the 3DPW dataset.
[…]
--- - --- - --- - --- - --- - 0.282 0.760 465801
$\checkmark$ --- - --- - --- - --- - 0.762 0.214 43824
$\checkmark$ $\checkmark$ --- - --- - --- - 0.839 0.146 25392
$\checkmark$ $\checkmark$ $\checkmark$ --- - --- - 0.882 0.129 12482
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ --- - 0.902 0.052 10647
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.921 0.028 10647
Ground Truth of 3DPW 0.937 --- - 7
Table 3: Evaluation of the proposed pipeline on web videos. PCK (threshold $0.1$) represents the consistency between the 2D pose and SMPL estimation, while SOIOU measures the proportion of the SMPL mask outside the SAM mask. #Sub. denotes the number of qualified video clips. P.X/S.X denotes the Protocol/Stage X.